# literature.bib

### v2

@@ 3.2 Non-Concentration of the Stationary Solution
where we abbreviated $e_{1}(t)$ for the first column of $E(t)$. Hence, the primal variable follows a Gaussian distribution at every time $t>0$, with mean $z_{0}^{T}e_{1}(t)$ and variance $\lVert(e_{1}(t)\rVert_{\Sigma_{0}}^{2}+2\int_{0}^{t}e^{2}_{11}(s)\,\mathrm{d}s$. Due to the contraction result Theorem 3.4, we know that $\mu_{t}$ converges to the stationary distribution $\mu^{\text{{stat}}}_{\text{{cPD}}}$ as $t\to\infty$. Indeed, this holds here since $\lim_{t\to\infty}E(t)=0$, so $\mathbb{E}_{X\sim\mu^{\text{{stat}}}_{\text{{cPD}}}}[X]=0$. The variance of $\mu^{\text{{stat}}}_{\text{{cPD}}}$ is
$\operatorname{Var}_{X\sim\mu^{\text{{stat}}}_{\text{{cPD}}}}[X]=2\int_{0}^{\infty}e^{2}_{11}(s)\,\mathrm{d}s=\frac{c_{g}(c_{f}+k^{2}c_{g})+\lambda c_{f}^{2}c_{g}^{2}}{(c_{f}+k^{2}c_{g})(1+\lambda c_{f}c_{g})},$
which does not coincide with the variance of the target distribution $\mu^{\ast}$. The target is however approximated as $\lambda$ becomes large, since $\lim_{\lambda\to\infty}\operatorname{Var}[\mu^{\text{{stat}}}_{\text{{cPD}}}]=c_{g}c_{f}/(c_{f}+k^{2}c_{g})$.
@@
$\operatorname{Var}_{Z\sim\pi^{\text{{stat}}}_{\text{{cPD}}}}[Z]=\frac{1}{(c_{f}+k^{2}c_{g})(1+\lambda c_{f}c_{g})}\begin{pmatrix}c_{g}(c_{f}+k^{2}c_{g})+\lambda c_{f}^{2}c_{g}^{2}&k\lambda c_{f}c_{g}^{2}\\ k\lambda c_{f}c_{g}^{2}&k^{2}\lambda c_{g}^{2}\end{pmatrix}$
@@
The negative result Theorem 3.5 naturally raises the question whether the diffusion equation can be modified such that the target becomes its stationary distribution. While we will consider the more general case of non-homogeneous, problem-dependent diffusion coefficients in Section 3.4, the (algorithmically) most straightforward approach would be to add homogeneous diffusion not only in the primal variable $X_{k}$ of the algorithm, but also in the dual $Y_{k}$. Unfortunately, the next theorem shows this approach can not be successful.
@@
Consider the same target $\mu^{\ast}$ as in Theorem 3.5, but a more general diffusion process
$\mathrm{d}Z_{t}=-A_{\text{{cPD}}}Z_{t}\mathrm{d}t+B_{\text{{cPD}}}\mathrm{d}W_{t},\quad\text{with }A_{\text{{cPD}}}=\begin{pmatrix}c_{g}^{-1}&k\\ -\lambda k&\lambda c_{f}\end{pmatrix},B_{\text{{cPD}}}=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\end{pmatrix}$ (3.8)
where $W_{t}$ is Brownian motion in $\mathbb{R}^{2}$. Then there exists no choice for the entries of $B_{\text{{cPD}}}\in\mathbb{R}^{2\times 2}$ independent of the constants $c_{g},c_{f}$ and $k$ such that the stationary solution $\pi^{\text{{stat}}}_{\text{{cPD}}}$ of (3.8) has $x$-marginal $\mu^{\ast}$.
@@
We start by noting that the contraction result Theorem 3.4 can easily be generalized to the more general $\Sigma_{\text{{cPD}}}=\frac{1}{2}B_{\text{{cPD}}}B_{\text{{cPD}}}^{T}$ with any constant coefficient $B_{\text{{cPD}}}$, since the relevant Lemma 3.3 holds for any constant diffusion coefficient. In view of this contraction property, we can assume the initialization $Z_{0}\sim\pi_{0}$ to be Gaussian with mean $z_{0}$, covariance $\Sigma_{0}$ and obtain $\pi^{\text{{stat}}}_{\text{{cPD}}}$ by computing $\lim_{t\to\infty}\pi_{t}$. The reason for initializing at a Gaussian is that we can again obtain the closed-form solution $\pi_{t}$ of the Cauchy problem for the Fokker-Planck equation. We carry out the same computations as in the proof of Theorem 3.5, showing that $\pi_{t}$ is Gaussian with mean $E(t)^{T}z_{0}$ and covariance matrix
$E(t)^{T}\Sigma_{0}E(t)+2\int_{0}^{t}E(s)^{T}B_{\text{{cPD}}}B_{\text{{cPD}}}^{T}E(s)\,\mathrm{d}s.$
The entries of $E$ can be computed explicitly (we omit the terms here) and by integrating and taking the limit $t\to\infty$, we find that the $x$-marginal of the stationary distribution $\pi^{\text{{stat}}}_{\text{{cPD}}}$ has variance
$\operatorname{Var}[\mu^{\text{{stat}}}_{\text{{cPD}}}]=\frac{(b_{11}^{2}+b_{12}^{2})(\lambda c_{f}^{2}c_{g}^{2}+c_{g}(c_{f}+k^{2}c_{g}))-2(b_{11}b_{21}+b_{12}b_{22})kc_{f}c_{g}^{2}+(b_{21}^{2}+b_{22}^{2})\lambda^{-1}k^{2}c_{g}^{2}}{2(1+\lambda c_{f}c_{g})(c_{f}+k^{2}c_{g})}.$
We want to show that there is no choice of $b_{ij}$ (independent of $c_{f},c_{g}$ and $k$) such that the target $\mu^{\ast}$ coincides with the $x$-marginal $\mu^{\text{{stat}}}_{\text{{cPD}}}$ of the stationary solution. The variance of both distributions is equal if and only if
@@
Isolating the terms including $k^{2}$ shows that this implies
@@
We want to elaborate a little more on the concentration properties of the stationary solution $\pi_{\text{{cPD}}}^{\text{{stat}}}$. Theorem 3.5 showed that in the limiting case $\lambda\to\infty$, the stationary distribution does not have a Lebesgue density anymore since the mass concentrates on the optimality condition $\partial f^{\ast}(y)=Kx$. In that case, the primal marginal actually converges to the target distribution, which is intuitively clear since plugging in $Y_{t}\in\partial f(KX_{t})$ into the equation for $X_{t}$ reduces the equation to overdamped Langevin diffusion. We will now generalize these properties of the toy example in Theorem 3.5 in two ways. Specifically, we will show in Section 3.3 under which conditions the stationary solution converges in the limit $\lambda\to\infty$ to a singular push-forward of the target $\mu^{\ast}$ to the space of joint distributions in $(x,y)$. Also, we now show that the concentration at the optimality generally does not occur for finite $\lambda$, even though the diffusion coefficient is degenerate. In the case of smooth coefficients, the Fokker-Planck operator can be shown to be hypoelliptic, implying that any solution to the dynamics has a smooth density with respect to the Lebesgue measure for all times $t>0$, even in the case of a concentrated initialization.
@@
The degenerate diffusion coefficient in (2.5) results in a constant but only positive semidefinite diffusion coefficient $\Sigma_{\text{{cPD}}}$ in the Fokker-Planck equation (3.7). In order to prove that the Fokker-Planck equation is hypoelliptic, we hence have to employ a weaker condition for hypoellipticity, usually termed ‘Hörmander’s condition’ [Hoermander1967]. In order to state it, we need the definition of the commutator operation and Lie algebras spanned by smooth vector fields.
@@
$\mathcal{L}_{\text{{cPD}}}u=-A_{\text{{cPD}}}^{T}(z)\nabla_{z}u+\Delta_{x}u.$
@@
Assume that $g\in C^{\infty}(\mathcal{X})$ and $f^{\ast}\in C^{\infty}(\mathcal{Y})$, so that the diffusion (2.6) has smooth coefficients. If $\dim(\mathcal{Y})=m\leq d=\dim(\mathcal{X})$ and $K$ has full rank $m$, then $\mathcal{L}_{\text{{cPD}}}$ satisfies Hörmander’s condition. As a consequence, the stationary solution $\pi^{\text{{stat}}}_{\text{{cPD}}}$ has a $C^{\infty}(\mathcal{Z})$-smooth Lebesgue density.
@@
We apply Lemma 3.9. $\mathcal{L}_{\text{{cPD}}}$ has the correct form with vector fields $A_{j}=\partial_{x_{j}}$, $j=1,\dots,d$ and
$A_{0}=-A_{\text{{cPD}}}^{T}(z)\nabla_{z}=\sum_{j=1}^{d}\left(-\frac{\partial g}{\partial x_{j}}(x)-k_{j}^{T}y\right)\partial_{x_{j}}+\lambda\sum_{i=1}^{m}\left((k^{i})^{T}x-\frac{\partial f^{\ast}}{\partial y_{i}}(y)\right)\partial_{y_{i}},$
@@ 3.3 Convergence to Overdamped Langevin Diffusion
We now continue to show that in the limiting case $\lambda\to\infty$ and under some smoothness assumption on $f$, the stationary solution $\pi^{\text{{stat}}}_{\text{{cPD}}}$ actually becomes singular, and further that the primal marginal distribution coincides in the limit with the target. Note that the formal case $\lambda=\infty$ was the subject of a recent paper [Habring2023], where the authors showed convergence in the case of perfect concentration $y\in\partial f(Kx)$.
Consider the rescaled model $(\lambda=\frac{1}{\varepsilon})$ taking the form
$\left\{\begin{array}[]{l}dX_{t}\in-(\partial g(X_{t})+K^{*}Y_{t})dt+\sqrt{2}dW_{t}\\ \varepsilon dY_{t}\in-(\partial f^{*}(Y_{t})-KX_{t})dt,\end{array}\right.$ (3.9)
where we again denote the distribution of $(X_{t},Y_{t})$ by $\pi_{t}$, with $x$-marginal $\mu_{t}$. We are now interested in the behaviour of $\mu_{t}$ in the limit $\varepsilon\rightarrow 0$. Intuitively, as $\varepsilon$ becomes small, in this formulation it becomes clear that $Y_{t}$ is forced to concentrate according to the optimality condition $Y_{t}\in\partial f(KX_{t})$. In this concentrated case, plugging in the dual variable into the primal inclusion shows that the dynamics actually coincide with overdamped Langevin dynamics, which has the correct invariant distribution. Time-discretizations of the limiting case $\varepsilon=0$ have been considered in [Habring2023], where an intermediate iterate $Y_{t}\in\partial f(KX_{t})$ is computed explicitly. Indeed, we can use a coupling argument between (3.9) and overdamped Langevin dynamics to bound the Wasserstein distance between the stationary state of (3.9) and the target.
@@
Assume that $f\in C^{3}$, that $g,f^{\ast}$ are $\omega_{g}$- respectively $\omega_{f^{\ast}}$-strongly convex, respectively, and that the target distribution $\mu^{\ast}$ satisfies
$\mathbb{E}_{X\sim\mu^{\ast}}\left[\lVert\nabla(\nabla f\circ K)(X)\rVert_{F}^{2}\right]=:C_{1}<\infty,$
$\mathbb{E}_{X\sim\mu^{\ast}}\left[\lVert\nabla\cdot\nabla(\nabla f\circ K)(X)\rVert^{2}\right]=:C_{2}<\infty,$
$\mathbb{E}_{X\sim\mu^{\ast}}\left[\lVert\nabla(\nabla f\circ K)(X)\left(\nabla g(X)+K^{T}\nabla f(KX)\right)\rVert^{2}\right]=:C_{3}<\infty,$
where in the second line the Laplacian is understood acting component-wise on the $\mathcal{Y}$-valued function $\nabla f\circ K$. For $0<\varepsilon<\omega_{f^{\ast}}/\omega_{g}$ and a constant $C(\varepsilon)=\sqrt{\varepsilon C_{1}/\omega_{g}}+\mathcal{O}(\varepsilon^{3/2})$, the following two statements hold:
@@
    If the primal-dual dynamics (3.9) are initialized with $X_{0}\sim\mu_{0}$ and $Y_{0}$ concentrated at the optimality condition $Y_{0}=\nabla f(KX_{0})$, then for every $t>0$ it holds
    $\mathcal{W}_{2}(\mu_{t},\mu^{\ast})\leq\left((1+\varepsilon\omega_{f^{\ast}}\lVert K\rVert)^{1/2}\mathcal{W}_{2}(\mu_{0},\mu^{\ast})+C(\varepsilon)\right)e^{-\omega_{g}t}+C(\varepsilon).$
@@
    The $x$-marginal $\mu^{\text{{stat}}}_{\text{{cPD}}}$ of the stationary solution $\pi^{\text{{stat}}}_{\text{{cPD}}}$ of (3.9) satisfies
    $\mathcal{W}_{2}(\mu^{\text{{stat}}}_{\text{{cPD}}},\mu^{\ast})\leq C(\varepsilon).$
@@
Assume $(X_{t},Y_{t})$ with distribution $\pi_{t}$ solves (3.9). We construct a coupling with a process $\tilde{X}_{t}$ which is initialized as $\tilde{X}_{0}\sim\mu^{\ast}$ and solves the usual overdamped Langevin dynamics
$d\tilde{X}_{t}\in-(\partial g(\tilde{X}_{t})+K^{T}\nabla f(K\tilde{X}_{t}))dt+\sqrt{2}dW_{t}$ (3.10)
with the same Brownian motion $W_{t}$ as in (3.9). Since $\mu^{\ast}$ is the invariant distribution of (3.10), it follows $\tilde{X}_{t}\sim\mu^{\ast}$ for all $t>0$. The joint measure $\Pi_{t}$ of $(X_{t},Y_{t},\tilde{X}_{t})$ solves the Fokker-Planck equation
$\frac{\partial}{\partial t}\Pi_{t}$ $=\nabla_{x}\cdot((p+K^{*}y)\Pi_{t})+\frac{1}{\varepsilon}\nabla_{y}\cdot((q-Kx)\Pi_{t})+\nabla_{\tilde{x}}\cdot((\tilde{p}+K^{T}\tilde{y})\Pi_{t})$ (3.11)
@@
where as before $p\in\partial g(x),q\in\partial f^{\ast}(y),\tilde{p}\in\partial g(\tilde{x})$. With the artificial, concentrated dual variable $\tilde{Y}_{t}=\nabla f(\tilde{X}_{t})$, we denote the singular joint law of $(\tilde{X}_{t},\tilde{Y}_{t})$ by $\pi^{\ast}=(I,\nabla f\circ K)_{\#}\mu^{\ast}$.
We now consider the transport distance $\mathcal{T}_{1,\varepsilon}$ on $\mathcal{P}(\mathcal{Z})$ in order to bound the distance between $\pi_{t}$ and $\pi^{\ast}$, which due to the concentration of $\pi^{\ast}$ can be written as
$\mathcal{T}_{1,\varepsilon}(\pi_{t},\pi^{\ast})=\inf_{\Pi\in\Gamma(\pi,\mu^{\ast})}\left(\mathbb{E}_{(x,y,\tilde{x})\sim\Pi}\left[d_{1,\varepsilon}(x,y,\tilde{x})^{2}\right]\right)^{1/2},$
@@
$d_{1,\varepsilon}(x,y,\tilde{x}):=\left(\lVert x-\tilde{x}\rVert^{2}+\varepsilon\lVert y-\nabla f(K\tilde{x})\rVert^{2}\right)^{1/2}.$
@@
$\frac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[d_{1,\varepsilon}(x,y,\tilde{x})^{2}\right]$
$=\frac{1}{2}\int d_{1,\varepsilon}(x,y,\tilde{x})^{2}\left(\nabla_{x}\cdot((p+K^{T}y)\Pi_{t})+\frac{1}{\varepsilon}\nabla_{y}\cdot((q-Kx)\Pi_{t})+\nabla_{\tilde{x}}\cdot((\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\Pi_{t})\right)\,\mathrm{d}(x,y,\tilde{x})$
$\qquad+\frac{1}{2}\int d_{1,\varepsilon}(x,y,\tilde{x})^{2}\left((\nabla_{x}+\nabla_{\tilde{x}})\cdot(\nabla_{x}+\nabla_{\tilde{x}})\Pi_{t}\right)\,\mathrm{d}(x,y,\tilde{x})$
@@
$\qquad-\varepsilon\int\left(M(\tilde{x})(\tilde{y}(\tilde{x})-y)\right)\cdot(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\,\mathrm{d}\Pi_{t}-\varepsilon\int\left(M(\tilde{x})(\tilde{y}(\tilde{x})-y)\right)\cdot(\nabla_{x}+\nabla_{\tilde{x}})\Pi_{t}d(x,y,\tilde{x})$

### v3

@@ 3.2 Non-Concentration of the Stationary Solution
where we abbreviated $e_{1}(t)$ for the first column of $E(t)$. Hence, the primal variable follows a Gaussian distribution at every time $t>0$, with mean $z_{0}^{T}e_{1}(t)$ and variance $\lVert(e_{1}(t)\rVert_{\Sigma_{0}}^{2}+2\int_{0}^{t}e^{2}_{11}(s)\,\mathrm{d}s$. Due to the contraction result Theorem 3.4, we know that $\mu_{t}$ converges to the stationary distribution $\mu_{\lambda}$ as $t\to\infty$. Indeed, this holds here since $\lim_{t\to\infty}E(t)=0$, so $\mathbb{E}_{X\sim\mu_{\lambda}}[X]=0$. The variance of $\mu_{\lambda}$ is
$\operatorname{Var}_{X\sim\mu_{\lambda}}[X]=2\int_{0}^{\infty}e^{2}_{11}(s)\,\mathrm{d}s=\frac{c_{g}(c_{f}+k^{2}c_{g})+\lambda c_{f}^{2}c_{g}^{2}}{(c_{f}+k^{2}c_{g})(1+\lambda c_{f}c_{g})},$
which does not coincide with the variance of the target distribution $\mu^{\ast}$. The target is however approximated as $\lambda$ becomes large, since $\lim_{\lambda\to\infty}\operatorname{Var}[\mu_{\lambda}]=c_{g}c_{f}/(c_{f}+k^{2}c_{g})$.
@@
$\operatorname{Var}_{Z\sim\pi_{\lambda}}[Z]=\frac{1}{(c_{f}+k^{2}c_{g})(1+\lambda c_{f}c_{g})}\begin{pmatrix}c_{g}(c_{f}+k^{2}c_{g})+\lambda c_{f}^{2}c_{g}^{2}&k\lambda c_{f}c_{g}^{2}\\ k\lambda c_{f}c_{g}^{2}&k^{2}\lambda c_{g}^{2}\end{pmatrix}$
@@
The negative result Theorem 3.5 naturally raises the question whether the diffusion equation can be modified such that the target becomes its stationary distribution. While we will consider the more general case of non-homogeneous, problem-dependent diffusion coefficients in Section 3.4, the (algorithmically) most straightforward approach would be to add homogeneous diffusion not only in the primal variable $X^{n}$ of the algorithm, but also in the dual $Y^{n}$, as in (2.4). Unfortunately, the next theorem shows that the time-continuous limit (2.8) always remains asymptotically inconsistent.
@@
Consider the same target $\mu^{\ast}$ as in Theorem 3.5, but the more general diffusion (2.8)
$\mathrm{d}Z_{t}=-A_{\text{{PD}}}Z_{t}\mathrm{d}t+B_{\text{{gPD}}}\mathrm{d}W_{t},\quad\text{with }A_{\text{{PD}}}=\begin{pmatrix}c_{g}^{-1}&k\\ -\lambda k&\lambda c_{f}\end{pmatrix},B_{\text{{gPD}}}=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\end{pmatrix}$ (3.9)
where $W_{t}$ is Brownian motion in $\mathbb{R}^{2}$. Then there exists no choice for the entries of $B_{\text{{gPD}}}\in\mathbb{R}^{2\times 2}$ independent of $c_{g},c_{f}$ and $k$ such that the stationary solution $\pi_{\lambda}$ of (3.9) has $x$-marginal $\mu^{\ast}$.
@@
We start by noting that the contraction result Theorem 3.4 can be applied also the more general $\Sigma_{\text{{gPD}}}=\frac{1}{2}B_{\text{{gPD}}}B_{\text{{gPD}}}^{T}$ with any constant coefficient $B_{\text{{gPD}}}$, since the relevant Lemma 3.3 holds for any constant diffusion coefficient. In view of this contraction property, we can assume the initialization $Z_{0}\sim\pi_{0}$ to be Gaussian with mean $z_{0}$, covariance $\Sigma_{0}$ and obtain $\pi_{\lambda}$ by computing $\lim_{t\to\infty}\pi_{t}$. The reason for initializing at a Gaussian is that we can again obtain the closed-form solution $\pi_{t}$ of the Cauchy problem for the Fokker-Planck equation. We carry out the same computations as in the proof of Theorem 3.5, showing that $\pi_{t}$ is Gaussian with mean $E(t)^{T}z_{0}$ and covariance matrix
$E(t)^{T}\Sigma_{0}E(t)+2\int_{0}^{t}E(s)^{T}B_{\text{{gPD}}}B_{\text{{gPD}}}^{T}E(s)\,\mathrm{d}s.$
The entries of $E$ can be computed explicitly (we omit the terms here) and by integrating and taking the limit $t\to\infty$, we find that the $x$-marginal of the stationary distribution $\pi_{\lambda}$ has variance
$\operatorname{Var}[\mu_{\lambda}]=\frac{(b_{11}^{2}+b_{12}^{2})(\lambda c_{f}^{2}c_{g}^{2}+c_{g}(c_{f}+k^{2}c_{g}))-2(b_{11}b_{21}+b_{12}b_{22})kc_{f}c_{g}^{2}+(b_{21}^{2}+b_{22}^{2})\lambda^{-1}k^{2}c_{g}^{2}}{2(1+\lambda c_{f}c_{g})(c_{f}+k^{2}c_{g})}.$
We want to show that there is no choice of $b_{ij}$ (independent of $c_{f},c_{g}$ and $k$) such that the target $\mu^{\ast}$ coincides with the $x$-marginal $\mu_{\lambda}$ of the stationary solution. The variance of both distributions is equal if and only if
@@
Isolating the terms containing $k^{2}$ shows that this implies
@@
We want to elaborate more on the concentration properties of $\pi_{\lambda}$ in the scheme (2.6) with primal diffusion. Theorem 3.5 showed that in the limiting case $\lambda\to\infty$, the stationary distribution does not have a Lebesgue density anymore since the mass concentrates on the optimality condition $\partial f^{\ast}(y)=Kx$. In that case, the primal marginal actually converges to the target distribution, which is intuitively clear since inserting $Y_{t}\in\partial f(KX_{t})$ into the equation for $X_{t}$ reduces the system to overdamped Langevin diffusion. We will rigorously generalize these properties of the toy example of Theorem 3.5 in two ways. Specifically, we show in Section 3.3 under which conditions the stationary solution converges, as $\lambda\to\infty$, to a singular push-forward of the target $\mu^{\ast}$ in the space of distributions over $\mathcal{Z}$. Also, we next show that the concentration at the optimal dual value generally does not occur for finite $\lambda$, even though the diffusion coefficient is degenerate.
In the case of smooth coefficients, the Fokker-Planck operator can be shown to be hypoelliptic, implying that any solution to the dynamics has a smooth density with respect to the Lebesgue measure for all times $t>0$, even in the case of a concentrated initialization.
@@
The degenerate diffusion coefficient in (2.6) results in a constant but only positive semidefinite diffusion coefficient $\Sigma_{\text{{PD}}}$ in the Fokker-Planck equation (3.8). In order to prove that the Fokker-Planck equation is hypoelliptic, we hence have to employ a weaker condition for hypoellipticity, usually termed ‘Hörmander’s condition’ [Hoermander1967]. In order to state it, we need the definition of the commutator operation and Lie algebras spanned by smooth vector fields.
@@
$\mathcal{L}_{\text{{PD}}}u=-A_{\text{{PD}}}^{T}(z)\nabla_{z}u+\Delta_{x}u.$
@@
Assume that $g\in C^{\infty}(\mathcal{X})$ and $f^{\ast}\in C^{\infty}(\mathcal{Y})$, so that the diffusion (2.6) has smooth coefficients. If $\dim(\mathcal{Y})=m\leq d=\dim(\mathcal{X})$ and $K$ has full rank $m$, then $\mathcal{L}_{\text{{PD}}}$ satisfies Hörmander’s condition. As a consequence, the stationary solution $\pi_{\lambda}$ has a $C^{\infty}(\mathcal{Z})$-smooth Lebesgue density.
@@
We apply Lemma 3.9. $\mathcal{L}_{\text{{PD}}}$ has the correct form with vector fields $A_{j}=\partial_{x_{j}}$, $j=1,\dots,d$ and
$A_{0}=-A_{\text{{PD}}}^{T}(z)\nabla_{z}=\sum_{j=1}^{d}\left(-\frac{\partial g}{\partial x_{j}}(x)-k_{j}^{T}y\right)\partial_{x_{j}}+\lambda\sum_{i=1}^{m}\left((k^{i})^{T}x-\frac{\partial f^{\ast}}{\partial y_{i}}(y)\right)\partial_{y_{i}},$
@@
While the smoothness assumption on $g,f^{\ast}$ is necessary for technical reasons, our numerical test validated that the dispersion in the dual variable is a more general phenomenon. Concentration in the dual variable can generally not to be expected in the primal-dual dynamics.
@@ 3.3 Convergence to Overdamped Langevin Diffusion
We now continue by showing that in the limiting case $\lambda\to\infty$ and under some smoothness assumption on $f$, the stationary solution $\pi_{\lambda}$ of (2.6) becomes singular. In the limit, the primal marginal distribution coincides with the target.
Consider a solution $(X_{t},Y_{t})$ of (2.6) with distribution $(X_{t},Y_{t})\sim\pi_{t}$, denoting the $x$-marginal by $X_{t}\sim\mu_{t}$. We are interested in the behaviour of $\mu_{t}$ in the limit $\lambda\rightarrow\infty$. Intuitively, as $\lambda$ becomes large, it is clear that $Y_{t}$ is forced to concentrate, with the (formally) limiting SDE taking the form
$dX_{t}$ $\in-(\partial g(X_{t})+K^{T}Y_{t}dt+\sqrt{2}dW_{t}$
$Y_{t}$ $\in\partial f(KX_{t})).$
This coincides with overdamped Langevin diffusion on the potential $g+f\circ K$, with the target $\mu^{\ast}$ as invariant distribution.
This suggests we can use a coupling argument between (3.3) and the primal-dual dynamics (2.6) to bound the Wasserstein distance between the stationary state of (2.6) and the target.
@@
Assume that $f\in C^{3}$, that $g,f^{\ast}$ are $\omega_{g}$- and $\omega_{f^{\ast}}$-strongly convex, respectively, and that the target distribution $\mu^{\ast}$ satisfies
$\mathbb{E}_{X\sim\mu^{\ast}}\left[\lVert\nabla(\nabla f\circ K)(X)\rVert_{F}^{2}\right]=:D_{1}<\infty,$
$\mathbb{E}_{X\sim\mu^{\ast}}\left[\lVert\nabla\cdot\nabla(\nabla f\circ K)(X)\rVert^{2}\right]=:D_{2}<\infty,$
$\mathbb{E}_{X\sim\mu^{\ast}}\left[\lVert\nabla(\nabla f\circ K)(X)\left(\nabla g(X)+K^{T}\nabla f(KX)\right)\rVert^{2}\right]=:D_{3}<\infty,$
with constants $D_{i}$, where in the second line the Laplacian is understood acting component-wise on the $\mathcal{Y}$-valued function $\nabla f\circ K$. For $\omega_{g}/\omega_{f^{\ast}}<\lambda$ and a constant $C_{1}(\lambda)=(\lambda\omega_{g}/D_{1})^{-1/2}+\mathcal{O}(\lambda^{-3/2})$ as $\lambda\to\infty$, the following two statements hold:
@@
    If the primal-dual dynamics (2.6) are initialized with $X_{0}\sim\mu_{0}$ for some $\mu_{0}\in\mathcal{P}_{2}(\mathcal{X})$ and $Y_{0}$ concentrated at the optimality condition $Y_{0}=\nabla f(KX_{0})$, then for every $t>0$ it holds
    $\mathcal{W}_{2}(\mu_{t},\mu^{\ast})\leq\left((1+\lambda^{-1}\omega_{f^{\ast}}\lVert K\rVert)^{1/2}\mathcal{W}_{2}(\mu_{0},\mu^{\ast})+C_{1}(\lambda)\right)e^{-\omega_{g}t}+C_{1}(\lambda).$
@@
    The $x$-marginal $\mu_{\lambda}$ of the stationary solution $\pi_{\lambda}$ of (2.6) satisfies
    $\mathcal{W}_{2}(\mu_{\lambda},\mu^{\ast})\leq C_{1}(\lambda).$
@@
Assume $(X_{t},Y_{t})$ with distribution $\pi_{t}$ solves (2.5). Now consider a coupling with a process $\tilde{X}_{t}$ which is initialized as $\tilde{X}_{0}\sim\mu^{\ast}$ and solves (3.3) with the same Brownian motion $W_{t}$ as in (2.5). Since $\mu^{\ast}$ is the invariant distribution of (3.3), it follows $\tilde{X}_{t}\sim\mu^{\ast}$ for all $t>0$. The joint measure $\Pi_{t}$ of $(X_{t},Y_{t},\tilde{X}_{t})$ solves the Fokker-Planck equation
$\frac{\partial}{\partial t}\Pi_{t}$ $=\nabla_{x}\cdot((p+K^{T}y)\Pi_{t})+\lambda\nabla_{y}\cdot((q-Kx)\Pi_{t})+\nabla_{\tilde{x}}\cdot((\tilde{p}+K^{T}\tilde{y})\Pi_{t})$ (3.10)
@@
where as before $p\in\partial g(x),q\in\partial f^{\ast}(y),\tilde{p}\in\partial g(\tilde{x})$. With the artificial, concentrated dual variable $\tilde{Y}_{t}=\nabla f(K\tilde{X}_{t})$, we denote the singular joint law of $(\tilde{X}_{t},\tilde{Y}_{t})$ by $\pi^{\ast}=(I,\nabla f\circ K)_{\#}\mu^{\ast}$.
We now consider the transport distance $\mathcal{T}_{1,\lambda^{-1}}$ on $\mathcal{P}(\mathcal{Z})$ in order to bound the distance between $\pi_{t}$ and $\pi^{\ast}$, which due to the concentration of $\pi^{\ast}$ can be written as
$\mathcal{T}_{1,\lambda^{-1}}(\pi_{t},\pi^{\ast})=\inf_{\Pi\in\Gamma(\pi,\mu^{\ast})}\left(\mathbb{E}_{(x,y,\tilde{x})\sim\Pi}\left[d_{1,\lambda^{-1}}(x,y,\tilde{x})^{2}\right]\right)^{1/2},$
@@
$d_{1,\lambda^{-1}}(x,y,\tilde{x}):=\left(\lVert x-\tilde{x}\rVert^{2}+\lVert y-\nabla f(K\tilde{x})\rVert_{\lambda^{-1}}^{2}\right)^{1/2}.$
@@
$\frac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[d_{1,\lambda^{-1}}(x,y,\tilde{x})^{2}\right]$
$=\frac{1}{2}\int d_{1,\lambda^{-1}}(x,y,\tilde{x})^{2}\left(\nabla_{x}\cdot((p+K^{T}y)\Pi_{t})+\lambda\nabla_{y}\cdot((q-Kx)\Pi_{t})+\nabla_{\tilde{x}}\cdot((\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\Pi_{t})\right)\mathrm{d}(x,y,\tilde{x})$
$\qquad+\frac{1}{2}\int d_{1,\lambda^{-1}}(x,y,\tilde{x})^{2}\left((\nabla_{x}+\nabla_{\tilde{x}})\cdot(\nabla_{x}+\nabla_{\tilde{x}})\Pi_{t}\right)\,\mathrm{d}(x,y,\tilde{x})$
@@
$\qquad-\frac{1}{\lambda}\int\left(M(\tilde{x})(\tilde{y}(\tilde{x})-y)\right)\cdot(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\,\mathrm{d}\Pi_{t}-\frac{1}{\lambda}\int\left(M(\tilde{x})(\tilde{y}(\tilde{x})-y)\right)\cdot(\nabla_{x}+\nabla_{\tilde{x}})\Pi_{t}d(x,y,\tilde{x})$